Johnson Electronics is considering extending trade credit to some customers previously considered poor risks. Sales would increase by $146,000 if credit is extended to these new customers. Of the new accounts receivable generated, 8 percent will prove to be uncollectible. Additional collection costs will be 6 percent of sales, and production and selling costs will be 72 percent of sales. The firm is in the 10 percent tax bracket. a. Compute the incremental income after taxes. Incremental income after taxes b. What will Johnsons incremental return on sales be if these new credit customers are accepted? (Input your answer as a percent rounded to 2 decimal places.) sales c. If the accounts receivable turnover ratio is 4 to 1, and no other asset buildup is needed to serve the new customers, what will Johnsons incremental return on new average investment be? (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.) Incremental return on new average investment
Answer #1


Solution: 146000 a) Incremental sales Less : Bad debts=146000*8% Less: collection cost=146000*6% Less: Selling cost me before